OpenStudy (anonymous):

2x+5y=19 Find the slope of the line

OpenStudy (experimentx):

-2/5

OpenStudy (anonymous):

the slope of the line is 2x+5y=19 5y=-2x+19 compare it with the formula y=mx+c we will get y =-2/5x+19/5 so m=-2/5 so the slope is -2/5